OpenStudy (allie2020):

Ryan buys a helmet for $20 and a bike that is on sale for $25 off . At the register the cash or takes an additional temper sent off the total purchase if you initial price of the bike is $260 what is the final amount Ryan will pay . 193.50 $207 $229 or $229.50

OpenStudy (allie2020):

Additional 10%

OpenStudy (allie2020):

Cashier'

OpenStudy (sylph):

Whoops, thought it was asking for how much you got off. \(260-25=x\) \((x+20)\times.10=y\) \((x+20)-y=your~answer\)